### v1

# Mind the Gap Why Neural Memory Fails Under Semantic Density

Abstract
The brain solves a problem that current AI architectures do not: storing specific episodic facts without corrupting general semantic knowledge. Neuroscience explains this capacity through Complementary Learning Systems theory, which posits that biological memory requires two distinct subsystems - a fast hippocampal system for episodic storage using sparse, pattern-separated representations, and a slow neocortical system for extracting statistical regularities through distributed, overlapping representations. Current AI systems lack this architectural separation, attempting to serve both functions through neural weights alone, and we argue that this approach faces fundamental limits.
We identify what we term the Stability Gap in online neural memory: fast-weight and associative mechanisms that write facts into shared continuous parameters at inference time collapse to near-random accuracy within tens of semantically related facts. Through the lens of semantic density ($\rho$), we show that this collapse occurs as few as $N=5$ facts at high density ($\rho>0.6$) or $N\approx 20$7575 at moderate density, a phenomenon we formalize as the Orthogonality Constraint. Critically, this failure is not about attention capacity or context length - it persists even with perfect attention and unlimited context, arising instead from write-time interference when storage and retrieval share the same continuous substrate.
[…]
We propose Knowledge Objects (KOs) as the missing hippocampal component: discrete, typed memory units with controlled vocabularies and explicit version chains that provide interference-free episodic storage. Paired with neural weights serving the neocortical role, Knowledge Objects enable a true complementary learning architecture, and our findings suggest that reliable AI memory may require this bicameral design.
Keywords.
neural memory, associative memory, complementary learning systems, episodic memory, retrieval-augmented generation, knowledge representation, transformers
1 Introduction
The brain maintains various distinct memory systems, including the hippocampus (which rapidly encodes specific episodes with minimal interference), and the neocortex (which can be thought to slowly extract statistical patterns across many experiences). This division is not an accident of biology, but a solution to a fundamental computational problem; the same representations cannot simultaneously support fast, precise episodic storage and slow, distributed semantic learning (McClelland et al., 1995). Current AI architectures lack this separation, attempting to handle both episodic facts and semantic knowledge within the same neural substrate - whether context windows, parametric weights, or fast-weight associative memory. We argue that for online episodic storage via inner-product-addressed superposition in shared continuous parameters, overlap-driven interference imposes fundamental limits unless the system introduces explicit pattern separation.
Defining Memory.
This paper distinguishes four forms of “memory” in AI systems that are often conflated in the literature. The first is context-window attention, where facts remain visible within the current input sequence; this form is volatile (lost when the session ends) and incurs cost that grows with the amount of remembered material. The second is slow parametric memory in model weights, where facts are compressed into parameters during pretraining or finetuning; this form offers vast capacity but is expensive to update and does not natively support surgical overwrite, deletion, or version semantics. The third is fast parametric memory via online or associative mechanisms, where new facts are written at inference time into shared continuous parameters; this form enables instant writes but proves susceptible to interference under semantic density. The fourth is discrete addressable storage, where facts are stored as explicit key-value records in external data structures; this form supports deterministic updates and avoids interference between entries entirely. Our analysis demonstrates that the first three forms cannot reliably support precise episodic memory under realistic conditions, and that the fourth—discrete addressing—is architecturally necessary for the hippocampal function.
The Stability Gap.
We identify a core failure mode that we term the Stability Gap: online neural memory collapses under semantic density. When facts share semantic overlap—for instance, “Paris is the capital of France” and “Lyon is a city in France”—their representations interfere during retrieval because their embedding vectors occupy nearby regions of the continuous space. We formalize this phenomenon as the Orthogonality Constraint and demonstrate empirically that fast-weight and associative mechanisms collapse to near-random accuracy within tens of semantically related facts, far below theoretical capacity bounds that assume orthogonal keys. This limitation is not a bug to be fixed with scale or architectural refinement; it is a geometric inevitability arising from inner-product retrieval in shared continuous storage, where the training objective actively encourages the semantic clustering that causes storage interference.
Production Failures.
Production systems have partially recognized this problem by adopting discrete external databases for cross-session memory—when users enable “memory” features in ChatGPT or Claude, their facts are stored in databases rather than encoded into model weights. However, current implementations suffer from two failure modes that explain why memory can appear unreliable even when retrieval accuracy appears adequate. Schema drift occurs when predicate naming is inconsistent across extractions, with our experiments showing only 40–70% schema match across models. Version ambiguity occurs when corrections fail to cleanly supersede prior values, with some models mentioning both old and new values rather than providing the canonical current state.
The Solution: A Bicameral Architecture.
We propose Knowledge Objects (KOs) as the missing hippocampal component in AI memory architecture. Knowledge Objects are discrete, typed memory units with controlled vocabularies that enforce schema consistency and explicit version chains that resolve ambiguity about current values. Paired with neural weights serving the neocortical role of slow semantic generalization, Knowledge Objects enable a true complementary learning architecture: fast, interference-free episodic storage operating alongside slow, distributed pattern extraction. A learned router classifies incoming queries and directs them to the appropriate subsystem based on whether they require precise factual retrieval or fuzzy semantic reasoning.
Contributions.
[…]
Scope and Non-Claims.
The Orthogonality Constraint applies to any system that attempts online storage of discrete facts in shared continuous parameters. It does not claim that LLMs cannot store facts; they obviously can, via slow pretraining. The constraint specifically concerns inference-time episodic memory. We do not claim that attention mechanisms are unnecessary, nor that larger models cannot memorize more facts during pretraining. Rather, we show that attention alone does not provide stable write-time memory under semantic interference—no regime exists in which increased attention capacity alone improves long-term factual retention under semantic density. Similarly, we do not claim that RAG is useless; we show that naive similarity-only RAG lacks the identity, schema, and version semantics required for true episodic memory, and that adding these capabilities effectively introduces KO-like structure.
Paper Structure.
Section 2 reviews related work. Section 3 presents the theoretical framework. Section 4 describes Knowledge Objects and the hybrid architecture. Section 5 presents experimental validation. Section 6 discusses implications.
Figure 1: The Memory Architecture Gap. (a) Pure neural approaches to online memory—context windows and fast weights—face the Orthogonality Constraint: interference grows with semantic density, causing collapse. Production systems have adopted discrete external storage (databases) but suffer from schema drift and version ambiguity. (b) Our hybrid architecture provides structured discrete storage (Knowledge Objects) with controlled vocabularies and version chains, connected to the LLM via a learned router. This explains why discrete storage is necessary and fixes the remaining failure modes.
2 Related Work
Associative Memory Capacity.
Classical work established fundamental limits on neural memory. Hopfield networks store approximately $0.14N$ patterns before “blackout catastrophe” (Hopfield, 1982; McEliece et al., 1987). Ramsauer et al. (Ramsauer et al., 2021) proved that transformer attention is equivalent to a modern Hopfield network update, inheriting these capacity constraints. Schlag et al. (Schlag et al., 2021) showed linear transformers enter an “overcapacity regime when sequence length exceeds key dimension,” establishing $O(d)$ bounds. Our work extends this analysis by introducing semantic density as a failure predictor: we show collapse occurs far below theoretical capacity when stored patterns share embedding similarity.
Episodic Memory for LLM Agents.
Recent work has argued that LLM agents require separate episodic memory modules to bridge parametric and context memory (Pink et al., 2025). While we share the conclusion that episodic memory separation is essential, our contribution is distinct: we provide mathematical and empirical evidence that such separation is necessary due to geometric interference limits (the Orthogonality Constraint), not merely beneficial for agent capabilities. The growing consensus around episodic memory modules validates our thesis; we supply the formal foundation explaining why this architecture emerges.
Stability Gap in Continual Learning.
The term “Stability Gap” has been used in continual learning to describe the drop in general capabilities when fine-tuning on new domains (Guo et al., 2024). Our usage is distinct: we define the Semantic Stability Gap as the precise threshold ($N$) at which episodic retrieval collapses to near-random accuracy due to interference in embedding space. The continual learning stability gap concerns capability degradation (forgetting how to perform general tasks); our stability gap concerns fact collision (inability to retrieve specific facts when semantically similar facts crowd the same vector space region).
Neural Episodic Memory.
Recent architectures like Titans (Munkhdalai et al., 2025) and Linear Transformers (Katharopoulos et al., 2020) learn “fast weights” at test time, treating memory as gradient optimization. While Titans reports 98.8% accuracy on associative recall, it provides no capacity bounds and does not characterize failure modes under semantic density. We show that the core mechanism underlying such architectures—superposition of key-value pairs in shared continuous parameters—leads to collapse when storing semantically related facts.
Memory-Augmented Neural Networks.
Neural Turing Machines (Graves et al., 2014) and Differentiable Neural Computers introduced external memory banks with soft attention-based addressing. Memory Networks (Weston et al., 2014) demonstrated multi-hop reasoning over explicit memory slots. These systems use continuous addressing (soft attention over memory locations), which we show suffers from the same interference problem as in-weight storage when keys are semantically similar. Our Knowledge Objects use discrete, hash-based addressing that guarantees zero interference.
Retrieval Augmented Generation.
RAG systems (Lewis et al., 2020) offload memory to external stores, retrieving from corpora at inference time. While modern vector databases support near-real-time upserts, RAG’s architecture is fundamentally append-first: new facts coexist with old ones, and there is no native mechanism to identify a canonical “current value” or cleanly supersede outdated information. Knowledge Objects differ in providing key-based identity for deterministic overwrite, controlled vocabularies for schema consistency, and version chains for explicit supersession semantics.
Graph-Based Knowledge Systems.
GraphRAG (Edge et al., 2024) constructs knowledge graphs from documents, enabling structured traversal. Our Knowledge Objects share the insight that typed, structured memory outperforms unstructured chunk retrieval. We differ in three ways: (1) key-based identity enabling deterministic overwrite rather than append-only indexing, (2) learned routing to classify query intent, and (3) version chains enabling temporal queries and clean supersession of outdated facts.
Complementary Learning Systems.
Our architecture is directly inspired by Complementary Learning Systems (CLS) theory from neuroscience (McClelland et al., 1995; Kumaran et al., 2016). CLS proposes that biological memory requires two systems with fundamentally different properties: a hippocampal system characterized by fast learning, sparse and discrete representations, and pattern separation that enables storage of specific episodes without interference; and a neocortical system characterized by slow learning, distributed and overlapping representations, and pattern completion that extracts statistical regularities through gradual exposure across many experiences. The hippocampus solves exactly the problem we address: storing arbitrary new facts quickly without corrupting existing memories, which it achieves through sparse, non-overlapping representations—the biological analog of discrete addressing. Our Knowledge Objects play the hippocampal role, while neural weights play the neocortical role, and sleep consolidation (which transfers hippocampal memories to cortex) maps to our proposed mechanism for promoting stable KO facts into model weights.
3 Theoretical Framework: The Orthogonality Constraint
Why does online neural memory collapse under semantic density? We analyze the interference dynamics of a stylized Linear Associative Memory (LAM) module, as a proxy for test-time “fast-weight” and associative memory mechanisms (e.g., Titans-style updates) that write new facts by superposition into a shared continuous matrix at inference time. Throughout this section, the Orthogonality Constraint refers to separation in the key/query embedding space used to address memory (dimension $d$, typically the hidden size), not to the total number of parameters of a pretrained LLM, nor to the ability of slow parametric weights to encode broad world knowledge after large-scale training.
Given $N$ stored fact pairs $\{(\mathbf{k}_{i},\mathbf{v}_{i})\}_{i=1}^{N}$, the memory matrix is:
$\mathbf{M}=\sum_{i=1}^{N}\mathbf{v}_{i}\otimes\mathbf{k}_{i}^{T}$ (1)
When retrieving fact $j$ by querying with $\mathbf{k}_{j}$:
$\mathbf{M}\cdot\mathbf{k}_{j}=\underbrace{\mathbf{v}_{j}(\mathbf{k}_{j}\cdot\mathbf{k}_{j})}_{\text{signal}}+\underbrace{\sum_{i\neq j}\mathbf{v}_{i}(\mathbf{k}_{i}\cdot\mathbf{k}_{j})}_{\text{interference}}$ (2)
The retrieval output is the sum of the intended signal ($\mathbf{v}_{j}$) plus an interference term from every other stored fact. Eq. (2) shows retrieval is the sum of a target signal and a multi-item interference term. Increasing $N$ under non-negligible semantic density $\rho$ causes interference to dominate, so the failure mode is stability/isolation under superposition (cross-talk), not a claim that pretrained weights cannot store large amounts of knowledge. If keys were orthogonal ($\mathbf{k}_{i}\cdot\mathbf{k}_{j}=0$ for $i\neq j$), interference would vanish. But real semantic embeddings have non-zero semantic density:
$\rho=\frac{1}{N(N-1)}\sum_{i\neq j}\cos(\mathbf{k}_{i},\mathbf{k}_{j})$ (3)
Retrieval fails when interference overwhelms signal. With $N$ facts at density $\rho$, the expected interference magnitude scales as $O(N\cdot\rho)$. This predicts: (1) collapse accelerates with $N$, and (2) higher $\rho$ causes earlier collapse. We note that this scaling argument is approximate: the tightest theoretical bound depends on both the mean and variance of pairwise cosines, and may scale as $O(\sqrt{N}\cdot\sigma_{\rho})$ under certain distributional assumptions. However, our measure $\rho$ (mean pairwise cosine) serves as an effective empirical predictor, and our experiments (Figure 5) validate the predictive relationship between $\rho$ and collapse threshold $N_{50}$ regardless of the precise theoretical bound.
Why High-Dimensional Capacity Bounds Don’t Help.
A natural objection is that high-dimensional spaces can accommodate exponentially many near-orthogonal vectors. Johnson-Lindenstrauss-style bounds suggest that for embedding dimension $d$ and maximum pairwise similarity $\varepsilon$, one can fit:
$N\lesssim\exp\left(\frac{\varepsilon^{2}\cdot d}{4}\right)$ (4)
For $d=7000$ (DeepSeek-scale) and $\varepsilon=0.1$, this permits $\sim 4\times 10^{7}$ vectors. Why then does collapse occur at $N\approx 20$?
The problem is that semantic embeddings cannot exploit this geometric capacity. The training objective for embedding models requires similar concepts to cluster together—this is what enables generalization and transfer. “Paris is the capital of France” and “Lyon is a city in France” must be nearby in embedding space because they share words, appear in similar contexts, and predict similar continuations.
The Orthogonality Constraint thus arises not from geometric limits on $d$, but from the fundamental tension between two objectives that cannot be simultaneously satisfied. The compression objective—what embeddings are trained for—requires grouping semantically similar content together to enable generalization and transfer learning. The storage objective—what factual recall requires—demands separating distinct facts to enable interference-free retrieval. These objectives are fundamentally opposed: the same geometric property (clustering similar concepts) that makes embeddings useful for prediction makes them unreliable for storage.
But Doesn’t Pretraining Solve This?
One might argue that pretraining already solves this problem: the model learns to “squeeze all possible semantic meanings into $d$-dimensional space” efficiently. If pretraining can pack exponentially many concepts, why can’t it pack facts?
The answer is that pretraining optimizes for prediction, not storage. These objectives are fundamentally opposed:
Prediction Storage
Goal Predict next token Retrieve exact fact
Similar inputs Should produce similar outputs Must remain distinguishable
“France facts” Should cluster (enables transfer) Must be separable (enables lookup)
Trained for Generalization Precision
The embedding model is trained to put “Paris is capital of France” near “Lyon is city in France” because they share words, appear in similar contexts, and predict similar continuations. This clustering is a feature for generalization—it enables the model to transfer knowledge about France to new contexts. But it is a bug for factual lookup—it causes “Paris” and “Lyon” to interfere with each other during retrieval.
If pretraining had successfully orthogonalized facts, we would not observe: (1) the “Lost in the Middle” phenomenon, (2) our measured stability gap collapse, (3) models hallucinating similar-but-wrong facts, or (4) semantic drift in long-context agents. But we observe all of these. The JL capacity exists in theory, but the training objective does not exploit it for fact storage—it exploits it for pattern compression.
Real-world knowledge exhibits $\rho\approx 0.3$$0.7$, not $\rho\approx 0$. This is why collapse occurs at $N\approx 20$—far below both the JL bound and the classical Hopfield capacity of $\sim 0.14d\approx 54$ for $d=384$.
Discrete Memory: True Orthogonality.
In discrete memory, keys are hash addresses; if $\mathbf{k}\neq\mathbf{k}_{\text{old}}$, overlap is exactly 0 by construction. The Stability Gap—the accuracy difference between discrete and neural memory—is thus a direct consequence of accumulated interference in continuous representations.
Figure 2: The Orthogonality Constraint. Left: In neural memory, semantically related facts (here, two statements about French cities) produce embedding vectors with small angular separation ($\cos\theta\approx 0.7$). When retrieving one fact, the system also partially retrieves others with similar keys—interference that grows with corpus size $N$ and semantic density $\rho$, causing collapse by $N\approx 20$7575 depending on density. Right: Discrete memory assigns unique hash addresses to each fact; overlap between addresses is exactly zero by construction. Retrieval returns $\mathbf{v}_{A}$ with no interference regardless of how many other facts share semantic similarity.
4 Methodology
This section presents our methodology in detail, beginning with the neural baseline that serves as our theoretical comparator, followed by the Knowledge Object architecture that constitutes our primary contribution, and concluding with the hybrid system that combines both approaches through learned routing.
4.1 Neural Memory Baseline: Linear Associative Memory
To represent gradient-based episodic memories such as Titans and Linear Transformers, we implement a Test-Time Training (TTT) Linear Memory, a simplified proxy that captures the essential dynamics of systems that learn via gradient updates to shared continuous weights. The storage mechanism employs a weight matrix $\mathbf{M}\in\mathbb{R}^{d\times d}$ to store all facts; given a key-value pair $(\mathbf{k},\mathbf{v})$, the matrix updates according to
$\mathbf{M}_{t+1}=\lambda\mathbf{M}_{t}+\eta(\mathbf{v}\otimes\mathbf{k}^{T})$ (5)
where $\lambda=0.99$ is a decay factor and $\eta=0.1$ is the learning rate. Retrieval proceeds by computing $\hat{\mathbf{v}}=\mathbf{M}\mathbf{q}$ for a query $\mathbf{q}$ and decoding via nearest-neighbor matching against stored values. This baseline captures the core failure mode we analyze: gradient updates to shared weights cause interference when keys overlap in the embedding space. Production systems may incorporate additional mechanisms such as normalization and regularization, but the fundamental constraint—that semantically similar keys produce blended retrievals—remains inherent to any architecture that stores facts via superposition.
4.2 Knowledge Objects: Online Episodic Memory
Knowledge Objects constitute the core methodological contribution of this work. Unlike RAG, which retrieves from static pre-indexed corpora, KOs provide online episodic memory: facts written during inference become available immediately, without requiring batch re-indexing. This capability addresses a fundamental limitation of RAG systems that has received insufficient attention in the literature.
4.2.1 The RAG Limitation
Consider a conversational agent in which a user states “My name is Alice.” Modern vector databases can perform near-real-time upserts, so the embedding could be added quickly. However, RAG’s architectural limitations run deeper than indexing latency. First, RAG is append-first: adding the new fact does not remove or supersede any prior statement about the user’s name, so both old and new values may be retrieved. Second, RAG lacks key-based identity: there is no canonical address for “the user’s name” that can be surgically updated; facts are identified only by embedding similarity, which is fuzzy. Third, RAG provides no schema guarantees: the fact might be stored as “user’s name is Alice” or “Alice is the user’s name” depending on phrasing, breaking downstream systems that expect consistent structure. Fourth, RAG has no native version semantics: if the user later says “Actually, call me Alicia,” the system has no mechanism to mark the prior value as superseded—both will be retrieved with similar scores.
These limitations are not addressed by standard similarity-only RAG; resolving them requires adding explicit identity, schema, and version semantics—effectively introducing KO-like structure. Knowledge Objects provide these capabilities natively: hash-based addressing provides key-based identity for surgical updates, controlled vocabularies enforce schema consistency, version chains provide explicit supersession semantics, and the combination enables true online episodic memory where facts are available immediately with deterministic overwrite behavior.
4.2.2 Knowledge Object Data Structure
Each Knowledge Object is represented as a typed tuple of the form
$\text{KO}=(\text{id},\text{subject},\text{predicate},\text{object},\text{embedding},\text{provenance})$ (6)
where the id is a unique identifier computed as $\text{hash}(\text{subject}\|\text{predicate})$, the subject denotes the entity about which the fact pertains (such as “user”, “patient_MR”, or “France”), the predicate specifies the attribute or relationship (such as “name”, “capital”, or “tumor_size”), and the object contains the value itself (such as “Alice”, “Paris”, or “3cm”). The embedding is a dense vector computed from the concatenation of subject, predicate, and object, enabling semantic search, while the provenance field captures metadata including source, timestamp, confidence score, and version pointer.
4.2.3 Storage: Hash-Addressed with $O(1)$ Writes
Knowledge Objects are stored in a hash map keyed by $\text{hash}(\text{subject}\|\text{predicate})$, which provides constant-time insertion and update, constant-time exact lookup when the subject and predicate are known, and zero interference between facts since distinct keys map to distinct addresses and updating one fact cannot corrupt another. This last property matters most: whereas neural memory systems store all facts in shared continuous parameters where updates necessarily interact, hash-based storage guarantees complete isolation between memory entries.
4.2.4 Retrieval: Semantic Search over Embeddings
For queries that do not specify an exact subject-predicate pair, we perform $k$-nearest neighbor search over the embedding space. The process begins by encoding the query as $\mathbf{q}=\text{Encoder}(\text{query})$, then finding the top-$k$ Knowledge Objects by cosine similarity via $\text{argmax}_{k}\cos(\mathbf{q},\mathbf{e}_{i})$, and finally returning the matched KOs to augment the LLM’s response. This hybrid approach—exact lookup when possible, semantic search otherwise—combines the precision of structured databases with the flexibility of neural retrieval.
4.2.5 The KO Compiler: Extracting Structured Facts
Raw text must be compiled into structured Knowledge Objects, a process we accomplish through an LLM-based compiler. The compiler receives natural language input and extracts factual statements as subject-predicate-object triples, using snake_case for predicates to ensure consistency and including confidence scores for each extraction. For instance, given the input “Dr. Sarah Chen is a 42-year-old cardiologist at MGH. Her phone is 617-555-0123,” the compiler produces four triples: (Dr_Sarah_Chen, age, 42), (Dr_Sarah_Chen, occupation, cardiologist), (Dr_Sarah_Chen, employer, MGH), and (Dr_Sarah_Chen, phone_number, 617-555-0123).
4.2.6 Surgical Updates: $O(1)$ Correction
When a fact changes, KO performs a surgical update by computing the identifier $\text{id}=\text{hash}(\text{subject}\|\text{predicate})$, optionally archiving the old value with a version pointer for history preservation, and writing the new value at the same address. This operation completes in $O(1)$ time regardless of corpus size. By contrast, RAG systems must append the new fact to a pending queue, re-index the entire corpus or affected partition, and handle the conflict period during which old and new facts coexist until re-indexing completes—a process that scales with corpus size and introduces temporal inconsistency.
4.2.7 Version Chains: Temporal Queries
Knowledge Objects optionally preserve version history, maintaining a linked list of prior values with timestamps. For example, a user’s favorite food might have evolved through versions: spaghetti (January 2024, archived), sushi (June 2024, archived), and pizza (January 2025, current). This structure enables temporal queries: asking “What is my favorite food?” returns the current value (pizza), asking “What was my favorite food before?” returns the previous version (sushi), and asking “What changed since January?” triggers traversal of the version chain. Neural memory fundamentally cannot support such queries because weight updates destroy prior states irrecoverably—there is no “undo” operation for gradient descent.
4.2.8 Key Properties Summary
Table 1 summarizes the architectural differences between Knowledge Objects and conventional RAG systems.
Table 1: Knowledge Objects vs RAG: Architectural comparison.
Property RAG Knowledge Objects
Fact identity Embedding similarity (fuzzy) Hash key (deterministic)
Update semantics Append (old persists) Supersede (clean overwrite)
Schema control None (free-form chunks) Controlled vocabulary
Data structure Unstructured chunks Typed (subject, predicate, object)
Version handling Both values retrieved Explicit version chains
Temporal queries Not supported Traverse version history
4.2.9 Illustrative Example: A Medical Research Assistant
To make the architecture concrete, consider a medical research assistant that helps a physician track patient cases over months of interaction. In a first session during January, the physician states: “Patient M.R., 67-year-old male, presented with persistent cough and 15-pound weight loss. CT scan showed 3cm mass in right upper lobe. Biopsy confirmed non-small cell lung cancer, stage IIB.” The KO compiler extracts structured facts—(patient_MR, age, 67), (patient_MR, diagnosis, NSCLC_IIB), (patient_MR, tumor_size, 3cm), and (patient_MR, tumor_location, right_upper_lobe)—which are written immediately to the KO store and become available for retrieval in subsequent queries without waiting for batch reindexing.
In a second session during March, the physician asks: “What was M.R.’s tumor size at diagnosis?” The router classifies this as a factual query, and KO retrieval returns 3cm deterministically via hash lookup on (patient_MR, tumor_size). The physician then reports: “Post-chemo scan shows tumor reduced to 1.2cm.” The system performs a surgical update in which the new value supersedes the prior one, but the version chain preserves history, enabling later queries such as “How has M.R.’s tumor responded to treatment?” that require temporal reasoning over the fact’s evolution.
Session 15 (November): The physician asks: “Summarize M.R.’s treatment trajectory.” The system retrieves the full version chain for patient_MR, constructs a temporal narrative, and generates: “M.R. presented in January with 3cm NSCLC in the right upper lobe. Following chemotherapy, tumor reduced to 1.2cm by March…”
This scenario illustrates the capabilities that distinguish Knowledge Objects from alternatives: facts become available immediately without waiting for reindexing (zero-latency writes), tumor size changes can be recorded without corrupting diagnosis or age information (surgical updates), queries about historical states can traverse the version chain (temporal queries), and factual queries hit the KO store while requests like “explain the treatment options” pass to the LLM (hybrid routing). A pure neural system would suffer interference as hundreds of patient facts accumulate, while a pure RAG system could not update tumor size without re-indexing the corpus.
4.3 Hybrid System: The Learned Router
Neither pure discrete memory nor pure neural generation handles all queries effectively. Factual queries such as “What is the capital of France?” benefit from KO’s interference-free retrieval, while fuzzy queries such as “Write a poem about Paris” should pass directly to the LLM. The hybrid architecture addresses this through a learned routing mechanism.
4.3.1 Architecture Overview
The system operates in four stages: first, the query is embedded using a sentence transformer; second, a learned router classifies the query as either “Factual” or “Fuzzy”; third, factual queries follow the fast path, retrieving from the KO store and augmenting the LLM context with the results; and fourth, fuzzy queries follow the slow path, passing directly to the LLM without KO retrieval.
4.3.2 Why Not Static Thresholds?
Initial experiments employed a static confidence threshold ($\tau=0.75$), where the system would use a KO match if its cosine similarity exceeded the threshold and fall back to the LLM otherwise. This approach failed under load because semantic crowding reduced cosine similarities below the threshold even for correct matches, triggering inappropriate fallback to the LLM when KO retrieval would have been appropriate.
4.3.3 The Learned Router
We experimented with two router variants: (1) logistic regression on raw query embeddings (384-dimensional vectors from the sentence transformer), and (2) logistic regression on handcrafted surface features extracted from the query text. Both approaches use a training set of 30 examples comprising 15 factual queries (such as “What is X?” and “Who created Y?”) and 15 fuzzy queries (such as “Write a poem” and “Explain the concept”). The surface-feature variant, which we report in our experiments, extracts features such as: presence of interrogative words (what, who, when, where), question mark termination, proper noun indicators, word count, and creative/explanatory verbs (write, help, suggest, explain). This variant achieved slightly better accuracy and interpretability, though both approaches exceeded 95% accuracy. We use logistic regression as the simplest router; in Experiment 5 (Table 6) we also evaluate other lightweight classifiers (Random Forest, SVM, Gradient Boosting) on the same feature set to show that the routing task is robustly learnable. What makes this work is that the router examines linguistic structure rather than embedding similarity scores, making it immune to vector space crowding: a query like “What is the capital of France?” is classified as factual regardless of how crowded the embedding space becomes.
Figure 3: Hybrid Architecture. Queries are embedded and classified by a Learned Router. Factual queries retrieve from the KO hash map; fuzzy queries pass directly to the LLM. This division of labor—discrete memory for facts, neural weights for generation—is the core architectural principle.
5 Experimental Results
All experiments used dense embeddings (all-MiniLM-L6-v2, $d=384$) and were run for 5 independent trials. We report mean $\pm$ standard deviation.
Definition of $N$.
Throughout, $N$ denotes the number of facts stored in memory. Each fact is a (key, value) pair where the key is an embedding and the value is the answer to retrieve.
Neural Baseline (LAM).
We implement Linearized Attention Memory as a proxy for neural memory systems that store facts via superposition in shared continuous parameters. Writing accumulates outer products: $\mathbf{W}\leftarrow\mathbf{W}+\mathbf{v}\otimes\mathbf{k}$ where $\mathbf{k}$ is the key embedding and $\mathbf{v}$ is the value embedding. Reading computes $\mathbf{r}=\mathbf{W}\mathbf{q}$ and finds the closest stored value. This simplified model captures the core interference dynamics: when keys are not orthogonal, their values blend during retrieval. Production systems may include additional mechanisms (gating, forgetting) that delay collapse, but cannot eliminate it when keys are semantically clustered.
Discrete Baseline (KO).
Knowledge Objects store facts as explicit key-value pairs in a hash map. Writing is $O(1)$ insertion. Reading supports both exact key lookup ($O(1)$) and top-$k$ semantic search ($O(N)$, or $O(\log N)$ with approximate nearest neighbor indexing). Version chains track history for each key.
5.1 Experiment 1: The Stability Gap
This experiment quantifies the fundamental accuracy gap between neural and discrete memory systems as the number of stored facts increases, validating the theoretical prediction that neural memory collapses under semantic density while discrete memory remains stable.
Methodology.
We evaluated both memory systems on synthetic entity-attribute-value facts following the template “What is the [attribute] of [entity]?” with answers drawn from a controlled vocabulary. The synthetic corpus was designed to exhibit high semantic density ($\rho\approx 0.70$), representing a challenging but realistic scenario where facts share similar linguistic structure—as would occur in a database of customer records, patient information, or product specifications where queries follow consistent patterns. We varied the corpus size $N$ from 1 to 75 facts in increments chosen to reveal the collapse trajectory. For each value of $N$, we stored all facts in both memory systems, then queried each fact and measured retrieval accuracy as the percentage of queries returning the correct answer. All experiments were run for 5 independent trials with different random seeds, and we report mean accuracy. The neural baseline (Linearized Attention Memory) was configured with embedding dimension $d=384$ using the all-MiniLM-L6-v2 sentence transformer, decay factor $\lambda=0.99$, and learning rate $\eta=0.1$. The discrete baseline (Knowledge Objects) used the same embedding model for semantic search, with facts stored in a hash map keyed by subject-predicate pairs.
Table 2: Stability Gap Results: Neural collapse under high semantic density ($\rho\approx 0.70$). Gap measured in percentage points ($p$).
N KO Acc. Neural Acc. Gap $\rho$
1 100% 100% 0pp 0.00
5 100% 40% 60pp 0.66
10 100% 20% 80pp 0.68
20 100% 5% 95pp 0.69
50 100% 4% 96pp 0.71
75 100% 1.3% 98.7pp 0.71
Analysis.
Neural memory achieves 100% accuracy at $N=1$, confirming correct implementation of the storage and retrieval mechanisms. As $N$ increases, accuracy collapses rapidly: 40% at $N=5$, 5% at $N=20$, and 1.3% at N=75N=75—barely above the random baseline of $1/75\approx 1.3\%$. Discrete memory maintains 100% accuracy throughout the entire range, yielding a stability gap of 98.7 percentage points at $N=75$. The high semantic density ($\rho\approx 0.70$) accelerates collapse compared to corpora with more diverse fact structures; this confirms that semantic density $\rho$ is a key predictor of neural memory failure, with higher density producing faster collapse.
Figure 4: The Stability Gap. Neural memory (red) collapses as NN increases, while discrete memory (green) maintains 100% accuracy throughout. Embedding dimension $d=384$ is held constant; only the number of stored facts varies. The shaded region represents the accuracy gap caused by write-time interference in continuous vector storage. Even as we store more facts, no improvement in attention or retrieval mechanism can prevent this collapse—it is a geometric consequence of semantic overlap in shared parameters.
Semantic Density Analysis.
To further validate the relationship between semantic density and collapse, we generated fact corpora with varying $\rho$ values by manipulating the diversity of entity names, attribute types, and question structures. Figure 5 plots $N_{50}$ (the number of facts at which neural accuracy drops to 50%) against semantic density. The relationship follows an exponential decay: at $\rho\approx 0.34$, collapse occurs around $N=20$; at $\rho>0.6$, collapse is nearly immediate with $N_{50}\leq 5$. We could not generate realistic fact corpora with $\rho<0.3$—even maximally diverse templates (different domains, question structures, entity types) produce baseline semantic similarity because natural language has inherent structure that embeddings capture.
Figure 5: Semantic Density Predicts Collapse. $N_{50}$ (the number of facts at which neural accuracy drops to 50%) decreases sharply as semantic density ($\rho$) increases. The exponential decay validates the Orthogonality Constraint: higher $\rho$ means more key overlap, hence more interference. Note: $\rho<0.3$ was not achievable with realistic fact structures.
5.2 Experiment 2: Reasoning Under Load
This experiment tests whether the interference observed in single-fact retrieval compounds when multi-hop reasoning is required, as would occur in question-answering tasks that require chaining multiple facts together.
Methodology.
We constructed a 3-hop reasoning chain requiring sequential fact retrieval: the query “What is the price of the main export of the capital of Zorbia?” requires first retrieving that Nexar is the capital of Zorbia, then that Crystite is the main export of Nexar, and finally that 500 credits is the price of Crystite. Success requires traversing all three hops correctly; an error at any hop propagates to the final answer. We tested both memory systems under two conditions: isolation, where only the 6 facts needed for the chain are stored ($N=6$), and under load, where 100 distractor facts with similar structure are added ($N=106$). The distractor facts follow the same entity-attribute-value template but concern unrelated entities, creating semantic interference without providing relevant information. We ran 5 independent trials for each condition.
Table 3: Reasoning Robustness: Multi-hop accuracy under semantic noise (5 trials).
System Isolation (N=6) Under Load (N=106)
KO (Discrete) 100.0% 100.0%
Neural (LAM) 100.0% 0.0%
Analysis.
Both systems succeed in isolation, demonstrating that neural memory can perform multi-hop reasoning when the vector space is uncontaminated by interfering facts. Under load, however, neural memory collapses completely to 0% accuracy while Knowledge Objects maintain perfect performance. The 100 distractor facts create sufficient semantic interference that even the first hop of the reasoning chain fails, preventing any possibility of reaching the correct final answer. This result has significant implications for practical applications: real-world knowledge bases contain thousands to millions of facts, and any system relying on neural memory for multi-hop reasoning will face catastrophic interference long before reaching useful scale.
5.3 Experiment 3: Deterministic Overwrite Semantics (KO vs RAG)
This experiment tests whether systems provide deterministic immediate overwrite semantics: when a fact is updated, can the system guarantee that (a) the new value is immediately available, and (b) the old value is cleanly superseded rather than persisting alongside the new one?
The RAG Limitation.
While modern vector databases support near-real-time upserts, RAG’s architectural limitations extend beyond indexing latency. RAG is fundamentally append-first and retrieval-first: new facts are added to the index, but there is no mechanism to identify and remove the prior value for the same logical fact. If a user says “My phone is 555-1111” and later “My phone changed to 555-2222,” both facts exist in the index with similar embeddings, and retrieval may return either or both depending on query phrasing and similarity scores. This is not a bug—it reflects RAG’s design for static corpora where facts accumulate rather than mutate.
Methodology.
We simulated a conversational session with 6 fact-learning exchanges, each following the same pattern: the user provides a fact in natural language (such as “My name is Alice” or “My birthday is March 15th”), the system extracts and stores the fact, and then we immediately query for that fact using a differently-phrased question (such as “What is my name?” or “When is my birthday?”). The key measurement is whether the system can retrieve the just-learned fact with deterministic accuracy. Our RAG baseline models a common production deployment pattern where new documents enter a pending queue for batch indexing—while some vector databases support near-real-time upserts, many production RAG systems use asynchronous ingestion pipelines for throughput and cost reasons. We also measured accuracy after a simulated batch reindex to establish an upper bound on eventual performance.
Table 4: Deterministic Overwrite: Can the system guarantee immediate, clean retrieval of updated facts? RAG baseline models batch-ingestion pipelines common in production deployments.
System Immediate Retrieval After Batch Reindex
KO (Discrete) 100.0% N/A
RAG (Batch-Ingestion) 0.0% 83.3%
Analysis.
Knowledge Objects retrieve newly-learned facts immediately with 100% accuracy because hash-based storage provides key-based identity: the fact “user’s name” has a deterministic address that is overwritten when updated. The batch-ingestion RAG baseline returns 0% before reindexing because new embeddings have not yet been added to the searchable index. Even after reindexing, RAG achieves only 83.3% accuracy due to embedding mismatch between the phrasing of statements (“My birthday is March 15th”) and queries (“When is my birthday?”), a known limitation of embedding-based retrieval when statement and query forms differ substantially. We note that RAG systems with real-time upsert capabilities would eliminate the “immediate retrieval” gap, but would still lack the key-based identity and version semantics needed for deterministic overwrite—both old and new values would coexist in the index with similar embeddings.
These results confirm that KO provides deterministic overwrite semantics that RAG architecturally lacks. Applications requiring conversational learning (where users teach the system facts and expect immediate recall), real-time personalization (where user preferences must be applied within the same session they are expressed), or surgical fact correction (where outdated information must be immediately superseded) require key-based identity and version semantics that RAG’s similarity-based retrieval cannot provide.
5.4 Experiment 4: Surgical Correction
This experiment tests whether facts can be updated individually without affecting other stored facts, a capability essential for maintaining accurate knowledge bases over time as information changes.
Methodology.
We wrote 100 facts to memory, then selected 10 facts at random and updated their values (for example, changing “Value_042” to “UPDATED_Value_042”). We then queried all 100 facts and measured accuracy separately for two groups: the 10 updated facts (which should return the new value) and the 90 unchanged facts (which should return their original values). This design tests both the effectiveness of updates (do updated facts reflect the new value?) and the isolation of updates (do other facts remain unaffected?). We ran 5 independent trials with different random selections of facts to update.
Table 5: Surgical Correction: Updating 10 facts in a 100-fact corpus (5 trials).
System Updated $\rightarrow$ New Value Others Preserved Operations
KO (Discrete) 100.0% 100.0% $O(1)$
Neural (LAM) 0.0% 1.1% $O(N)$
Analysis.
Knowledge Objects achieve perfect surgical updates: 100% of updated facts return the new value, and 100% of other facts remain unchanged, with each update completing in $O(1)$ time regardless of corpus size. Neural memory exhibits catastrophic forgetting: 0% of updates succeed (the system returns neither old nor new values reliably), and 98.9% of other facts are corrupted by the update operation. This occurs because gradient-based updates to shared weight matrices necessarily affect all stored associations, not just the target fact. Discrete addressing provides complete isolation between facts—updating one cannot affect another—while neural memory fundamentally cannot support surgical correction without catastrophic interference.
5.5 Experiment 5: Mixed Workload (Factual + Fuzzy)
Real-world workloads mix factual queries that seek specific retrievable answers (such as “What is the capital of France?”) with fuzzy queries that require generation or reasoning (such as “Write a poem about Paris”). A practical memory system must route each query type appropriately: factual queries should retrieve from the discrete KO store, while fuzzy queries should pass directly to the LLM. We tested whether a learned router can correctly classify query intent with minimal training data.
Methodology.
We constructed a test corpus of 100 queries, comprising 50 factual queries (who/what/when/where questions with specific retrievable answers) and 50 fuzzy queries (creative, open-ended, or explanatory requests). The router is a logistic regression classifier trained on only 15 labeled examples from each category. The feature vector captures linguistic patterns that distinguish factual from fuzzy queries: whether the query contains interrogative words (what, who, when, where, which), whether it ends with a question mark, whether it contains proper nouns indicated by capitalization, the total word count, and whether it contains creative or explanatory verbs (write, help, suggest, explain, describe). These features were chosen to capture the structural differences between factual lookup requests and open-ended generation requests without requiring semantic understanding.
Table 6: Mixed Workload: Router classification accuracy (10-fold CV).
Classifier Accuracy (95% CI)
Random Forest (100 trees) 97.8% [95.0%, 100%]
SVM (Linear) 96.5% [93.2%, 99.8%]
Logistic Regression 95.3% [90.5%, 100%]
Gradient Boosting 95.3% [90.3%, 100%]
Analysis.
Even simple classifiers achieve 95%+ routing accuracy with 10-fold cross-validation, demonstrating that the factual/fuzzy distinction is learnable from surface features without deep semantic analysis. Feature ablation reveals that the most informative features are the presence of temporal question words (removing “has_when” drops accuracy by 4.6 percentage points) and definiteness markers (“is the” and “my” each contribute 3.5 percentage points). The learning curve analysis shows that 50 labeled examples suffice to reach 90%+ accuracy, making the router practical to deploy with minimal human annotation effort. We also tested adversarial queries designed to confuse the classifier, such as “What’s a creative way to remember the capital of France?” which blends factual content with creative framing; the router achieves 83% accuracy on these edge cases, demonstrating reasonable robustness while identifying an area for future improvement.
5.6 Experiment 6: Cost Analysis
Beyond accuracy, we analyze the economic viability of context-based versus selective retrieval, a consideration that determines whether a memory architecture can be deployed at enterprise scale.
Methodology.
We measured actual token counts for queries at varying corpus sizes by constructing fact corpora of N=1, [^0] N=1{,}000, 5,0005{,}000, and 10,00010{,}000 facts and counting the tokens required to represent them. Context-based memory must load all $N$ facts into the prompt for every query because the LLM needs visibility into the entire knowledge base to locate the relevant fact. With facts averaging approximately 13 tokens each (in our synthetic corpus), the input token count per query is approximately $100+(N\times 13)+20$ tokens, comprising the system prompt, all facts, and the query. At N=5, [^0] N=5{,}000, this yields approximately 65,000 input tokens per query.
Knowledge Objects, by contrast, use embedding-based retrieval to identify the top-$k$ relevant facts before calling the LLM, so only retrieved facts enter the prompt. With $k=10$, the input token count per query is approximately $100+(10\times 13)+20\approx 250$ tokens, which remains constant regardless of corpus size $N$. The retrieval step itself (embedding lookup plus approximate nearest neighbor search) costs less than \$0.0001 per query and adds negligible latency.
We computed monthly costs using Claude Sonnet 4 pricing (\$3 per million input tokens, \$15 per million output tokens), assuming 50 output tokens per response and query volumes representative of realistic deployments.
Table 7: Monthly cost comparison: Context loads all $N$ facts per query; KO retrieves top-10 ($k=10$). Token counts measured empirically.
N Tokens/Query Queries/day Context Cost KO Cost
1,000 13,000 100 \$120/mo \$4/mo
5,000 65,000 500 \$2,940/mo \$19/mo
10,000 130,000 1,000 \$11,760/mo \$38/mo
Analysis.
[…]
5.7 Experiment 7: Production Memory Failure Modes
This experiment moves beyond synthetic benchmarks to evaluate how production memory systems actually fail in realistic use cases, testing frontier model memory capabilities against four potential failure modes that users might encounter.
Methodology.
We designed four tests targeting distinct failure modes. All API calls used temperature $t=0.0$ and top_p =1.0=1.0 to maximize reproducibility; we note that schema drift persists even at zero temperature, indicating it arises from the model’s learned priors rather than sampling stochasticity. For Claude models, we used the Anthropic API with model identifiers claude-sonnet-4-5-20250929; for GPT models, we used the OpenAI API with identifiers gpt-5.2 and gpt-4o-mini. The extraction prompt was: “Extract facts from the statement as JSON. Use this exact format: {“entity”: “…”, “predicate”: “…”, “value”: “…”}. Respond with ONLY the JSON, no other text.” No function calling or structured output mode was used; we tested the models’ natural extraction behavior.
For schema consistency, we provided 10 facts in natural language and asked the model to extract them as structured subject-predicate-object triples, measuring whether the model used consistent predicate names across extractions (for example, always using birth_place rather than varying between birth_place, was_born_in, and place_of_birth). For correction handling, we provided a sequence of statements where facts were updated (for example, “Alice’s phone is 555-1111” followed later by “Alice’s phone changed to 555-2222”), then queried for the current value, measuring whether the model returned only the latest value or ambiguously mentioned both old and new values. For retrieval accuracy, we loaded 20 facts into context and queried each, measuring simple retrieval performance. For positional bias, we varied the position of a target fact within a 50-fact context to test whether retrieval accuracy depended on where facts appeared in the prompt.
Table 8: Production memory failure mode analysis (Claude Sonnet 4.5).
Test Claude KO (Structured)
Schema consistency 70% 100% (controlled vocab)
Correction handling 0% 100% (version chains)
Retrieval (N=20) 100% 100%
Positional bias (N=50) 100% 100% (hash lookup)
Schema Drift.
When extracting structured facts, Claude Sonnet 4.5 matched our expected predicate names only 70% of the time. The 30% mismatches included variations such as birth_place being extracted as was_born_in, founded_year becoming founded_in, and floor becoming is_on_floor. While these variations are semantically sensible, they break downstream systems that rely on exact predicate matching for lookup. A database query for WHERE predicate = ’birth_place’ will miss facts stored under was_born_in, causing silent retrieval failures that users experience as the system “forgetting” facts it actually stored. Knowledge Objects address this through a controlled vocabulary enforced during extraction: the KO compiler constrains predicates to a pre-defined ontology, ensuring consistent naming regardless of how users phrase their inputs. We note that function-calling and structured output modes can mitigate schema drift, and we view this as evidence for our thesis: production-grade memory requires a constrained interface with controlled vocabularies, not free-form generative extraction.
Version Ambiguity.
When facts were corrected over the course of a conversation, Claude Sonnet 4.5 returned only the latest value 0% of the time—in all cases, it mentioned both old and new values in its response (for example, “Alice’s phone was 555-1111, but it changed to 555-2222”). While this behavior might seem helpful by providing context, it creates ambiguity for downstream systems and users who need definitive current values. If a user asks “What is Alice’s phone number?” they typically want the current number, not a history lesson. Knowledge Objects address this through explicit version chains that ensure corrections supersede prior values: queries return only the current version by default, with history available only when explicitly requested.
Retrieval and Positional Bias.
At moderate scale ($N\leq 100$), all models achieved perfect retrieval accuracy with no detectable positional bias—facts at the beginning, middle, and end of context were retrieved equally well. This confirms that the Stability Gap we observed in Experiment 1 affects neural memory (gradient-based storage in weights) rather than context-window attention at small $N$. The attention mechanism successfully retrieves facts from context when $N$ is small; the problems arise from schema drift and version ambiguity, not retrieval failure.
Multi-Model Comparison.
We extended our analysis to GPT-5.2 and GPT-4o-mini to verify that these findings generalize across frontier models from different providers rather than reflecting idiosyncrasies of Claude’s implementation. We tested schema consistency under two conditions: multi-session, where each of 10 facts was extracted in a separate API call (simulating realistic production use where facts arrive over hours, days, or weeks), and single-session, where all 10 facts were extracted in a single API call (allowing the model to observe its own prior extractions within the same context and potentially maintain consistency by reference).
Table 9: Multi-model comparison: Schema drift persists even within single sessions. Corrections column shows percentage of queries returning only the latest value without mentioning old values.
Model Schema (Multi) Schema (Single) Corrections Retrieval Cost/mo
Claude Sonnet 4.5 70% 70% 0% 100% \$5,642
GPT-5.2 60% 70% 100% 100% \$4,700
GPT-4o-mini 50% 60% 100% 100% \$282
KO (Proposed) 100% 100% 100% 100% \$28
Key Finding: Single-Session Provides Minimal Improvement.
[…]
Implications.
The failures we observe are not specific to any single model but reflect architectural limitations shared across frontier models from both Anthropic and OpenAI. All tested models exhibit schema drift, with each model producing different variants: birth_place becomes was_born_in (Claude), place_of_birth (GPT-5.2), or born_in (GPT-4o-mini). Interestingly, the GPT models handle corrections cleanly (100% returning only the latest value) while Claude consistently produces ambiguous responses mentioning both old and new values (0% clean corrections). This divergence suggests different internal approaches to representing and reasoning about temporal information across model families, though neither approach achieves the schema consistency required for reliable structured storage. Knowledge Objects address both problems through controlled vocabularies that constrain predicate choices during extraction and explicit version chains that ensure corrections cleanly supersede prior values without requiring the model to “remember” to omit old information.
5.8 Summary: Claude Memory vs Knowledge Objects
Table 10 synthesizes our findings into a direct comparison between Claude’s production memory system and Knowledge Objects, using Claude Sonnet 4.5 results from the multi-model comparison.
Table 10: Production comparison: Claude Memory vs Knowledge Objects at $N=5,000$ facts. Claude results from Sonnet 4.5 (multi-model comparison).
Metric Claude Memory Knowledge Objects
Retrieval accuracy 100% 100%
Schema consistency 70% 100%
Correction handling 0% 100%
Monthly cost (500 q/day) \$5,642 \$28
Latency per query 1.7s 0.2s
Cost scaling $O(N)$ $O(1)$
The Brute-Force Workaround.
[…]
Knowledge Objects match Claude’s retrieval accuracy while providing four additional guarantees that production systems require. First, schema guarantees through controlled vocabularies prevent predicate drift by constraining extraction to a pre-defined ontology, ensuring that “phone number” is always stored as phone_number regardless of how the user phrases it. Second, version guarantees through explicit chains ensure that corrections cleanly supersede prior values, so queries return only the current version unless the user explicitly requests history. Third, economic viability through $O(1)$ cost per query enables deployment at enterprise scale where context-based approaches would be prohibitively expensive. Fourth, interactive latency through sub-second response times supports real-time applications where users expect immediate feedback.
6 Discussion
The Core Finding.
Neural memory systems that store facts in shared continuous parameters will experience interference proportional to semantic density, and only systems employing discrete addresses can guarantee zero interference between stored facts. This finding does not represent a limitation of current models that might be resolved through scaling or architectural refinements—rather, it reflects a geometric constraint inherent to continuous vector storage that no amount of engineering can circumvent. When two embedding vectors occupy nearby regions of the space, any linear combination that retrieves one will partially retrieve the other; this is not a bug but a mathematical property of inner products.
Why Biology Got There First.
The brain confronted this problem millions of years ago: how to rapidly learn specific episodic facts (“that particular berry made me sick yesterday”) without corrupting general semantic knowledge (“berries are food”). Evolution’s solution, formalized in cognitive neuroscience as Complementary Learning Systems theory, mirrors our architecture with striking precision. The hippocampus employs sparse, pattern-separated representations that enable fast, interference-free storage of specific episodes, while the neocortex uses distributed, overlapping representations optimized for slow extraction of statistical regularities across many experiences. Sleep consolidation gradually transfers stable hippocampal memories to cortical storage, allowing the hippocampus to clear space for new episodes.
Current AI systems attempt to accomplish everything within the “neocortex”—that is, neural weights and attention mechanisms—which explains why they fail at precise factual storage despite succeeding at pattern recognition, generation, and reasoning. Our Knowledge Objects implement the missing “hippocampus” in this analogy: not as a biological simulation, but as the engineering solution to the same computational problem that evolution solved through anatomical specialization.
Why This Matters for RAG.
Retrieval-Augmented Generation addresses the interference problem by offloading storage to external databases, but introduces limitations that become critical for dynamic knowledge bases. While modern vector databases support near-real-time upserts, RAG’s architecture is fundamentally append-first and retrieval-first: new facts are added alongside old ones, and there is no native mechanism to identify the canonical “current value” for a logical fact or to cleanly supersede outdated information. For static corpora such as legal archives or technical documentation, these limitations are acceptable because the knowledge base rarely mutates. For conversational agents, personal assistants, or any system that must learn and update during deployment, the lack of key-based identity and version semantics creates the schema drift and version ambiguity failures we document in Experiment 7.
Knowledge Objects address these limitations: hash-based addressing provides key-based identity for deterministic lookup and surgical updates, controlled vocabularies enforce schema consistency during extraction, and version chains provide explicit supersession semantics where corrections cleanly replace prior values.
The Learned Router.
Static similarity thresholds fail under semantic load because vector space crowding reduces cosine similarities below threshold even for correct matches, triggering inappropriate fallback to generative responses when retrieval would have been appropriate. The learned router succeeds by classifying query structure rather than relying on similarity scores. A Random Forest classifier trained on standard linguistic features achieves 97.8% routing accuracy (95% CI: [95.0%, 100%]), with our learning curve analysis demonstrating that 50 labeled examples suffice to reach 90%+ accuracy. This means the router can be deployed with minimal human annotation effort and adapted to new domains with a small labeling investment.
Relation to Production Systems.
Production AI assistants have empirically converged on discrete external storage for cross-session memory, validating the necessity of our approach. When users enable “memory” features in commercial systems, facts are stored in external databases rather than encoded into model weights. However, current implementations typically store unstructured text chunks such as “User mentioned they like Python” rather than structured triples such as (user, preferred_language, Python). This distinction matters significantly: structured storage enables deterministic lookup by subject and predicate, surgical updates that change single facts without affecting others, and version chains that track fact evolution over time. Our multi-model comparison demonstrates that schema drift and version ambiguity affect both Claude and GPT model families, confirming that these are architectural limitations of the unstructured approach rather than bugs in any particular model’s implementation.
Limitations.
Our analysis has several limitations that should inform interpretation of the results. First, our neural baseline (LAM) is a simplified proxy for production systems that store facts via superposition; we treat it as a mechanistic lower bound that isolates the interference mechanism arising whenever semantically overlapping keys are stored in shared continuous parameters. Production implementations may include mitigations such as gating and regularization that shift the collapse curve rightward, but absent explicit pattern separation, overlap-driven interference persists when keys are semantically clustered—our claim concerns the existence of interference under semantic density, not the performance of any specific implementation. Second, our experiments use a single embedding model (all-MiniLM-L6-v2 with $d=384$ dimensions); different embedding dimensions would shift the quantitative results, though the qualitative finding—that collapse occurs at $N$ far below theoretical capacity—should hold. Third, we measure retrieval accuracy on synthetic fact corpora, not end-to-end task performance on downstream applications, though retrieval accuracy is a prerequisite for task success. Fourth, our multi-model comparison tests only three models from two providers; additional models might reveal different failure patterns, though the consistency we observe suggests these are general phenomena.
7 Conclusion
This paper has characterized three fundamental problems with context-based AI memory that together explain why such systems can appear unreliable despite their impressive capabilities on other tasks.
The first problem concerns theoretical limits on storage capacity. The Stability Gap we documented shows that neural memory collapses to near-random accuracy as semantic density increases, while discrete memory maintains perfect retrieval regardless of corpus size. This gap arises from what we term the Orthogonality Constraint: when embedding keys share semantic overlap, interference during retrieval becomes mathematically inevitable in continuous vector space. Absent mechanisms that enforce pattern separation (such as discrete identity, explicit orthogonalization at write time, or other anti-cross-talk structure), overlap-driven interference is unavoidable in inner-product-addressed superposition memory.
The second problem concerns correctness failures in production systems. We found that production memory systems fail not from retrieval degradation—which is handled adequately at moderate scale—but from schema drift (inconsistent predicate naming during fact extraction) and version ambiguity (old values persisting alongside new ones after correction). Our multi-model comparison demonstrates that these failures affect all tested frontier models, including Claude Sonnet 4.5, GPT-5.2, and GPT-4o-mini, with schema drift persisting even within single sessions where models can observe their own prior extractions. This confirms that the failures reflect architectural limitations inherent to generative extraction rather than bugs in any particular model’s implementation.
[…]
We introduced Knowledge Objects as a structured memory system addressing all three problems through complementary mechanisms: discrete hash-based storage prevents interference between facts, controlled vocabularies prevent schema drift during extraction, explicit version chains prevent ambiguity when facts are corrected, and selective retrieval achieves $O(1)$ cost per query regardless of corpus size.
Our findings establish that reliable, economical AI memory requires a hybrid architecture that divides labor according to computational strengths: neural weights for reasoning, pattern recognition, and generalization; discrete storage for precise factual recall; and a learned router to direct each query to the appropriate subsystem. The path to trustworthy AI memory runs not through ever-larger context windows or ever-denser weight matrices, but through principled separation of concerns that respects the fundamental capabilities and limitations of each storage mechanism.
Future Work.
Several directions merit further investigation. First, scaling discrete memory to millions or billions of facts will require integration with approximate nearest neighbor search techniques that maintain sub-linear query time. Second, trust-weighted retrieval that accounts for source reliability, recency, and corroboration could improve accuracy on contested or time-sensitive facts. Third, tighter integration with production LLMs could reduce hallucination rates on factual queries by conditioning generation on retrieved Knowledge Objects. Finally, the relationship between the Orthogonality Constraint and other forms of interference in neural networks (such as catastrophic forgetting during continual learning) deserves theoretical investigation.
Acknowledgments
We thank Peter Norvig, Suhair Khan, Ravin Kumar, and Anthony Rose for helpful discussions on earlier drafts.
Appendix A Knowledge Object Specification
This appendix provides the complete specification for Knowledge Objects, including the data structure, compiler prompts, and operational semantics.
A.1 Data Structure
A Knowledge Object is a 6-tuple:
KnowledgeObject {
    id: string           // hash(subject || predicate)
    subject: string      // The entity (e.g., "patient_MR", "France")
    predicate: string    // The attribute (e.g., "capital", "tumor_size")
    object: any          // The value (e.g., "Paris", 3.0)
    embedding: float[d]  // Dense vector, d=384 for MiniLM
    provenance: {
        source: string       // Origin of the fact
        timestamp: datetime  // When extracted
        confidence: float    // Extraction confidence [0,1]
        version: int         // Version number
        previous: string?    // Pointer to prior version (optional)
A.2 KO Compiler Prompt
The following prompt template is used to extract structured facts from unstructured text:
SYSTEM: You are a knowledge extraction system. Extract all
factual statements as structured triples.
RULES:
1. Subject: The primary entity the fact is about
2. Predicate: The relationship or attribute (use snake_case)
3. Object: The value, related entity, or state
4. Extract EVERY concrete fact, no matter how small
5. Include confidence score (0.0-1.0) for each extraction
6. For numerical values, preserve units
7. For temporal facts, include timestamps when available
OUTPUT FORMAT (JSON):
  {"subject": "...", "predicate": "...", "object": "...",
   "confidence": 0.95},
USER: [Input text to extract from]
A.3 Operational Semantics
Write(text).
Given input text:
  1. 1.
    Pass text to KO Compiler, receive list of triples
  2. 2.
    For each triple $(s,p,o)$:
    1. (a)
      Compute $\text{id}=\text{hash}(s\|p)$
    2. (b)
      Compute $\text{embedding}=\text{Encoder}(s+\text{`` ''}+p+\text{`` ''}+o)$
    3. (c)
      If id exists in store: archive current version, increment version number
    4. (d)
      Write KO to store at address id
  3. 3.
    Return list of created/updated KO ids
Read(query).
Given a natural language query:
  1. 1.
    Compute $\mathbf{q}=\text{Encoder}(\text{query})$
  2. 2.
    Find top-$k$ KOs by $\cos(\mathbf{q},\mathbf{e}_{i})$
  3. 3.
    Return matched KOs with scores
ReadExact(subject, predicate).
Given exact keys:
  1. 1.
    Compute $\text{id}=\text{hash}(\text{subject}\|\text{predicate})$
  2. 2.
    Return KO at address id (or null if not found)
Update(subject, predicate, new_object).
Surgical correction:
  1. 1.
    Compute $\text{id}=\text{hash}(\text{subject}\|\text{predicate})$
  2. 2.
    Archive current KO with version pointer
  3. 3.
    Create new KO with incremented version, previous pointer
  4. 4.
    Write to store at same address id
A.4 Implementation Notes
Hash Function.
We use SHA-256 truncated to 64 bits for the id hash. Collision probability is negligible for corpus sizes below $10^{9}$ facts.
Embedding Model.
Default: all-MiniLM-L6-v2 ($d=384$). For production, consider text-embedding-3-large ($d=3072$) for higher semantic resolution.
Index Structure.
The KO store maintains two indices: a hash index that maps id $\rightarrow$ KO for $O(1)$ exact lookup, and a vector index (HNSW or IVF) for approximate nearest neighbor search over embeddings.
Versioning.
Version chains are implemented as linked lists. Each KO stores a pointer to its previous version. Traversing the chain enables temporal queries.